how to derive unit of pole strength as Am show me the derivation of units for pole strength

Grade:12

1 Answers

rajneesh chander
6 years ago
we know, torque=MXB
M = torqe/ B
= Nm/Tesla 
= Nm/(N/m/A)     remember Tesla = force/q x vel.  (lorentz mag. force)
=Am2
now M=m x 2l
m= M/2l
= Am2/m
=Am